James is five years older than Thomas. Three years ago, James was twice as old as Thomas was. How old is James now?

Answer by macston(5194)

.
T=Thomas' age; J=James' age=T+5yrs
.
J-3yrs=2(T-3yrs)
(T+5yrs)-3yrs=2T-6yrs
T+2yrs=2T-6yrs
8yrs=T
Thomas is 8 years old.
.
J=T+5yrs=8yrs+5yrs=13yrs
ANSWER: James is 13 years old.